OpenStudy (anonymous):

solve equation by using the square root . Express radicals in simpliest form (3k-2)^2 = 16

OpenStudy (anonymous):

Two get rid of a square (^) you sqrt both sides.. Any time you sqrt both sides you need to put +- on the side without the x. So if you sqrt both sides.. you get 3k - 2 = +-4 You actually have two problems 3k -2 = 4 or 3x - 2 = -4 Then you solve each one by adding 2 and dividin by 3 3k = 6.... k = 3 3x = -2... k = -2/3 So k = -2/3 or 3

OpenStudy (anonymous):

solution set is {-2/3,3}

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

thank you